Show that. The curvature and torsion of the straight line is zero.
1
Expert's answer
2020-12-13T18:54:45-0500

Consider straight line: "y=cx+d"

Curvature of a twice differentiable function "y=f(x)" is given by "k=\frac{|y''|}{(1+(y')^2)^\frac{3}{2}}"

"y'=(cx+d)'=c"

"y''=(c)'=0"

"k=\frac{|0|}{(1+c^2)^\frac{3}{2}}=0"

Let "r(t)=\{x_0+p_1t,y_0+p_2t,z_0+p_3 t\}"  be the parametric equation of straight line. Then we can find the torsion using following formula:

"k_2=\frac{x'''(y'z''-y''z')+y'''(x''z'-x'z'')+z'''(x'y''-x''y')}{(y'z''-y''z')^2+(x''z'-x'z'')^2+(x'y''-x''y')^2}"

then:

"k_2=\frac{0\cdot(p_2\cdot0-0\cdot p_3)+0\cdot(0\cdot p_3-p_1\cdot 0)+0\cdot(p_1\cdot 0-0\cdot p_2)}{(p_2\cdot0-0\cdot p_3)^2+(0\cdot p_3-p_1\cdot 0)^2+p_1\cdot 0-0\cdot p_2')^2}= \frac{0}{0}" i.e. undefined